Date:
Friday, September 11, 2015
Telescope:
CTIO Blanco 4-m
Submitter:
James Lasker
Submitter Email:
Observer Support:
Manuel Hernandez
Night Assistant:
Alberto Alvarez
Problems:
SISPI crashed in the middle of flats multiple times. Did not process/transfer many VR/r flats the first time (exposures 473734-747) . Delete/ignore any files in that range that got through,we reran the VR and r flats after restarting SISPI
Incredibly high ellipticity around 04:15 CLT due to wind
Program:
Proposal:
2012B-0001
Target of Oportunity:
No
Time Observed:
5.0
Time Engineering:
0.0
Time lost technical:
0.0
Time lost weather:
0.0
Time lost other:
0.0
Description:
Dark Energy Survey
Primary Investigator:
Josh Frieman
Email:
Institution:
Fermilab
Instrument:
CTIO 4m DECam
Organization:
NOAO
Total Program Time:
5
Proposal:
2015B-0601
Target of Oportunity:
No
Time Observed:
5.0
Time Engineering:
0.0
Time lost technical:
0.0
Time lost weather:
0.0
Time lost other:
0.0
Description:
DECam Search for the Elusive Stellar Counterpart of the Magellanic Stream
Primary Investigator:
J. Chaname
Email:
Institution:
Pontificia Universidad Catolica de Chile
Instrument:
CTIO 4m DECam
Organization:
Chile
Total Program Time:
5
Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
---|---|---|---|---|---|---|
2015-09-11 | 10.0 | 0.0 | 0.0 | 0.0 | 0.0 | 10 |
*Observing Team*
_DES Observers:_ Ross Cawthon, James Lasker, Josh Frieman
_Telops:_ Manuel Hernandez, Alberto Alvarez
_Remote Support:_ ---
*Observing Plan*
#Standard Bias and Dome Flat sequence (See notes to DM)
# evening Standard stars
# Start ObsTac (Expecting gri band until SN fields up)
# Hand off to Community Observers at 11:15 CLT
# Take over from Community Observers at 4:15 CLT
# Restart ObsTac
# morning standards
*Conditions*
# Afternoon: Clear and sunny. No forecast for clouds at all.
# All night was clear.
# Seeling was poor in first quarter of night (1.4"), but got better in 4th quarter of night(as low as 0.8"). Small jump in psf due to observing into high wind.
*Accomplishments*
# Afternoon: zeros done, (ugrizYVR) dome flats done(See notes to DM), checked, OK, GPS Monitor Skyplot checked
# Evening standards SDSSJ1442-0005, E6-A (Skipped E8-A due to time constraints)
# aTmCam checked
# Early night : Obstac started, begins with z and Y band wide field exposures. Seeing ~1.2"
# At 20:50 CLT, Obstac tests the waters with a g and r band exposure. Finds 1.6" seeing... Wisely jumps back to z and Y.
# At 21:15 CLT, Obstac decides that it wants to do g and r band wide field so much that it will tolerate 1.5" seeing. Seeing settles down to 1.3" by 21:40 CLT (Some i and Y band sneaks in with predictably better seeing)
# Obstac alternates between g/r during good seeing and z/Y during bad seeing
# Finish off the first quarter night with SN-E1.
# Handoff to Community Observers at 23:15 CLT (At some point, wind screens closed and remain so)
# Takeover from Community Observers at 4:15 CLT
# Start with SN-S2 (see notes for DM), then SN-C3-z , C3-i, and C2
# Closed Louvers at 5:08 CLT (in the middle of SN-C3-z) due to wind above 17 mph
# Finish the night with two z band wide field exposures.
#Morning standards SDSSJ0100-0000, SDSSJ0200-0000, SDSSJ0320-0000
*Minor Problems and Relevant Information*
#
*Notes for Data Management*
# Exposures 473734-747 were taken surrounding a SISPI crash. Many did not get transferred, but please mark those that did a failed.
# Exposures 473980-85 (SN-S2) have very high ellipticity values, likely due to poinjting directly into high winds
# Possible small amount of occlusion on exposures 473999-474000 (Donuts checked by eye and they weren't conclusive, but looked similar to donuts from earlier in the day)
# During Exposure 474001 the dome rotated since we passed through Zenith